1/1

William Blake,

The Sick Rose

O Rose thou art sick. The invisible worm, That flies in the night In the howling storm:

Has found out thy bed Of crimson joy: And his dark secret love Does thy life destroy.

Verkinto de tiu ĉi Angla poemo estas William Blake (∗1757-11-28 – †1827-08-12).

Arg-974-1961 (2014-03-13 02:17:35)

Pri la poeto vidi la vikipediejon http: // eo. wikipedia. org/ wiki/ William_Blake .

http://www.poezio.net/version?poem-id=974&version-id=1961